# Information Loss and Cost in Symmetry Breaking

### v1

@@ Abstract
We develop an algebraic and information–theoretic framework to characterize symmetry breaking of generalized, non-invertible symmetries in two spatial dimensions. The reduction of symmetry is modeled within the subfactor theory, where condensable Frobenius algebras play the role of subgroups in the categorical setting. This perspective naturally connects to the description of anyon condensation in topological phases of matter. Central to our approach are coarse-graining maps, or conditional expectations, which act as quantum channels projecting observables from a phase with higher symmetry onto one where the symmetry is partially or completely broken by condensation. By employing relative entropy as an entropic order parameter, we quantify the information loss induced by condensation and establish a universal bound governed by the Jones index, equal to the quantum dimension of the condensate. We illustrate the framework through explicit examples, including the toric code, abelian groups $\mathbb{Z}_{N}$, and the representation category $\mathrm{Rep}(S_{3})$, and show how dualities give rise to equivalence classes of condensation patterns. Our results forge new connections between operator algebras, tensor category theory, and quantum information in the study of generalized symmetries.
@@ 1 Introduction
Related ideas have been presented in [26, 27, 28], although our work provides a comprehensive treatment that combines concepts from condensed matter, operator algebras, tensor category theory, and quantum information in the study of non-invertible symmetries. From condensed matter, we use the concept of anyon condensation [29]. In topologically ordered systems, anyon condensation refers to the process by which a subset of anyons—quasiparticle excitations in two-dimensional systems—condenses, giving rise to a new phase with modified topological order. Within the language of modular tensor categories, this phenomenon is captured by condensable algebras, a natural encoder of a symmetry sector in the case of non-invertible symmetries. These algebras identify the subset of bosonic anyons that can condense consistently with the ambient topological structure, thus encoding the symmetry-breaking patterns of the resulting phases [30].
@@
The work is organized as follows. Section 2 introduces the algebraic preliminaries, reviewing the structure of von Neumann algebras, quantum states, and the definition of relative entropy, which serves as a key information-theoretic tool throughout the work. In Section 3 we present subfactor theory in the language of anyon condensation, interpreting symmetry-breaking transitions as inclusions of algebras $\mathcal{N}\subset\mathcal{M}$ and explaining the role of Frobenius algebras in encoding symmetry-broken sectors. Section 4 constructs the conditional expectation as a quantum channel that implements the restriction map from the uncondensed to the condensed phase and introduces the lifting map that reverses this process. In Section 5, the relative entropy between the original and lifted states is proposed as an entropic order parameter, and the framework is illustrated through detailed examples including the groups $\mathbb{Z}_{N}$, the toric code, and the representation category of $S_{3}$. We also include an appendix discussing how the presence of dualities in the original theory implies dual condensation patterns leading to equivalent symmetry-broken patterns.
@@ 2 Background on von Neumann Algebras and states
$\omega(\alpha a+\beta b)=\alpha\,\omega(a)+\beta\,\omega(b)\,,\hskip 14.22636pt\omega(aa^{\dagger})\geq 0\,,\hskip 14.22636pt\omega(\mathbf{1})=\mathbf{1}\,.$ (2.5)
@@ 3 Subfactor Theory, Anyon Condensation and Entropic Order parameters
that is, $\mathcal{M}_{+}$ consists of all positive semidefinite operators in $\mathcal{M}$. If no such a conditional expectation exist, the index is defined to be $\mathcal{M}:\mathcal{N}]=\infty$. The index $\lambda=[\mathcal{M}:\mathcal{N}]$ quantifies the relative "size" of $\mathcal{M}$ compared to $\mathcal{N}$, analogous to the index of a subgroup but enriched by the analytic structure of operator algebras. It also provides an upper bound on the dimension of $\mathcal{M}$ when viewed as a module over $\mathcal{N}$
@@
In this work, since we are interested in the patterns of symmetry breaking in theories with generalized symmetries, we focus solely on determining the structure of $\mathcal{T}$, given the original algebra $\mathcal{A}$ and basic information about the condensate, such as the identification of the bosonic particles that condense. In this sense, we do not address the full topological phase transition from $\mathcal{A}$ to $\mathcal{U}$, but instead concentrate solely on the symmetry breaking from $\mathcal{A}$ to $\mathcal{T}$.  [^1] We note that the mathematical literature regarding anyon condensation uses category theory and the Frobenius algebra formalism and $Q$-systems [45]. A $Q$-system is a mathematical object that encodes a $C^{*}$-Frobenius algebra in a fusion category consisting of a triple $(\mathbb{A},\mu,\Delta)$ where: $\mathbb{A}$ is an object in a fusion category (for example, an MTC), $\mu:\mathbb{A}\otimes\mathbb{A}\to\mathbb{A}$ is a multiplication morphism and $\Delta:\mathbb{A}\to\mathbb{A}\otimes\mathbb{A}$ is a co-multiplication morphism. These morphisms satisfy associativity, co-associativity, and Frobenius conditions ensuring $\mathbb{A}$ forms a Frobenius algebra. In anyon condensation, $\mathbb{A}$ represents the condensable algebra. Therefore, if an algebra $\mathcal{A}$ corresponding to MTC describes the uncondensed phase, the condensation process leads to a subalgebra $\mathcal{T}$, induced by the $Q$-system..
@@ 3.1 Conditional expectation for the restriction and lifting maps
The trace-preserving property of the conditional expectation $\varepsilon_{\mathbb{A}}$ can be verified as follows:
@@ Acknowledgements
Appendix A Dualities and equivalent condensations
@@
$\mathbb{K}^{\mathbb{A}}_{a}$ $=\sum_{t\in\mathcal{T}}\,\lambda_{ta}^{\mathbb{A}}\,\Lambda_{ta}\,,\quad\lambda_{ta}^{\mathbb{A}}=\sqrt{(n^{t}_{a})_{\mathbb{A}}\left(\frac{d_{t}}{d_{a}}\right)}\,$ (A.3)
@@
$\varepsilon_{\mathbb{A}}(\mathrm{Q}\,\rho\,\mathrm{Q}^{\dagger})\equiv\varepsilon_{\mathbb{B}}(\rho)$ (A.4)
@@
$\lambda_{ta}^{\mathbb{B}}=\sum_{b\in\mathcal{A}}\,q_{ab}\,\lambda_{tb}^{\mathbb{A}}$ (A.5)
@@
$\lambda_{\phi 1}^{\mathbb{A}}=\lambda_{\phi Y}^{\mathbb{A}}=1=\lambda_{\mathbb{X}X}^{\mathbb{A}}=\lambda_{\mathbb{X}Z}^{\mathbb{A}}=1\,$, (A.6)
@@
$\varepsilon_{\mathbb{A}}(\rho)$ $=p_{\phi}|\phi\rangle\langle\phi|+p_{\mathbb{X}}|\mathbb{X}\rangle\langle\mathbb{X}|$ (A.7)
@@
$\lambda_{\phi 1}^{\mathbb{B}}=\lambda_{\phi Z}^{\mathbb{B}}=1=\lambda_{\mathbb{X}X}^{\mathbb{B}}=\lambda_{\mathbb{X}Y}^{\mathbb{B}}=1\,$, (A.8)
@@
$\varepsilon_{\mathbb{B}}(\rho)$ $=p_{\phi}|\phi\rangle\langle\phi|+p_{\mathbb{X}}|\mathbb{X}\rangle\langle\mathbb{X}|$ (A.9)
@@
${\rm Q}=\left(\begin{array}[]{cc}\mathbb{I}_{2}&0\\ 0&X\end{array}\right)\,$, (A.10)
@@
$X=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\,$, (A.11)
@@
$\varepsilon_{\mathbb{A}}(\mathrm{Q}\,\rho\,\mathrm{Q}^{\dagger})\equiv\varepsilon_{\mathbb{B}}(\rho)\,.$ (A.12)

### v2

@@ Abstract
We develop an algebraic and information–theoretic framework to characterize symmetry breaking of generalized, non-invertible symmetries in two spatial dimensions. The reduction of symmetry is modeled within the subfactor theory, where condensable Frobenius algebras play the role of subgroups in the categorical setting. This perspective naturally connects to the description of anyon condensation in topological phases of matter. Central to our approach are coarse-graining maps, or conditional expectations, which act as quantum channels projecting observables from a phase with higher symmetry onto one where the symmetry is partially or completely broken by condensation. By employing relative entropy as an entropic order parameter, we quantify the information loss induced by condensation and establish a universal bound governed by the Jones index, which is equal to the quantum dimension of the condensate. We illustrate the framework through explicit examples, including the toric code, abelian groups $\mathbb{Z}_{N}$, and the representation category $\mathrm{Rep}(S_{3})$, and show how dualities give rise to equivalence classes of condensation patterns. Our results forge new connections between operator algebras, tensor category theory, and quantum information in the study of generalized symmetries.
@@ 1 Introduction
Related ideas have been presented in [26, 27, 28], although our work provides a comprehensive treatment that combines concepts from condensed matter theory, theory of operator algebras, tensor category theory, and quantum information theory in the study of non-invertible symmetries. In condensed matter theory, we use the concept of anyon condensation [29]. In topologically ordered systems, anyon condensation refers to the process by which a subset of anyons—quasiparticle excitations in two-dimensional systems—condenses, giving rise to a new phase with modified topological order. Within the language of modular tensor categories, this phenomenon is captured by condensable algebras, a natural encoder of a symmetry sector in the case of non-invertible symmetries. These algebras identify the subset of bosonic anyons that can condense consistently with the ambient topological structure, thus encoding the symmetry-breaking patterns of the resulting phases [30].
@@
The work is organized as follows. Section 2 introduces the algebraic preliminaries, reviewing the structure of von Neumann algebras, quantum states, and the definition of relative entropy, which serves as a key information-theoretic tool throughout the work. In Section 3 we present subfactor theory in the language of anyon condensation, interpreting symmetry-breaking transitions as inclusions of algebras $\mathcal{N}\subset\mathcal{M}$ and explaining the role of Frobenius algebras in encoding symmetry-broken sectors. Section 4 constructs the conditional expectation as a quantum channel that implements the restriction map from the uncondensed to the condensed phase and introduces the lifting map that reverses this process. In Section 5, the relative entropy between the original and lifted states is proposed as an entropic order parameter, and the framework is illustrated through detailed examples including the groups $\mathbb{Z}_{N}$, the toric code, and the representation category of $S_{3}$. We also include two appendices. In Appendix A we prove some properties of the conditional expectation of Section 3 and Appendix B discusses how the presence of dualities in the original theory implies dual condensation patterns leading to equivalent symmetry-broken patterns.
@@ 2 Background on von Neumann Algebras and states
$\omega(\alpha a+\beta b)=\alpha\,\omega(a)+\beta\,\omega(b)\,,\hskip 14.22636pt\omega(aa^{\dagger})\geq 0\,,\hskip 14.22636pt\omega(\mathbf{1})=1\,.$ (2.5)
@@ 3 Subfactor Theory, Anyon Condensation and Entropic Order parameters
that is, $\mathcal{M}_{+}$ consists of all positive semidefinite operators in $\mathcal{M}$. If no such a conditional expectation exists, the index is defined to be $\mathcal{M}:\mathcal{N}]=\infty$. The index $\lambda=[\mathcal{M}:\mathcal{N}]$ quantifies the relative "size" of $\mathcal{M}$ compared to $\mathcal{N}$, analogous to the index of a subgroup but enriched by the analytic structure of operator algebras. It also provides an upper bound on the dimension of $\mathcal{M}$ when viewed as a module over $\mathcal{N}$
@@
In this work, since we are interested in the patterns of symmetry breaking in theories with generalized symmetries, we focus solely on determining the structure of $\mathcal{T}$, given the original algebra $\mathcal{A}$ and basic information about the condensate, such as the identification of the bosonic particles that condense. In this sense, we do not address the full topological phase transition from $\mathcal{A}$ to $\mathcal{U}$, but instead concentrate solely on the symmetry breaking from $\mathcal{A}$ to $\mathcal{T}$.  [^1] We note that the mathematical literature regarding anyon condensation uses category theory and the Frobenius algebra formalism and $Q$-systems [45]. A $Q$-system is a mathematical object that encodes a $C^{*}$-Frobenius algebra in a fusion category consisting of a triple $(\mathbb{A},\mu,\Delta)$ where: $\mathbb{A}$ is an object in a fusion category (for example, an MTC), $\mu:\mathbb{A}\otimes\mathbb{A}\to\mathbb{A}$ is a multiplication morphism and $\Delta:\mathbb{A}\to\mathbb{A}\otimes\mathbb{A}$ is a co-multiplication morphism. These morphisms satisfy associativity, co-associativity, and Frobenius conditions ensuring $\mathbb{A}$ forms a Frobenius algebra. In the theory of anyon condensation, $\mathbb{A}$ represents the condensable algebra. Therefore, if an algebra $\mathcal{A}$ corresponding to MTC describes the uncondensed phase, the condensation process leads to a subalgebra $\mathcal{T}$, induced by the $Q$-system..
@@ 3.1 Conditional expectation for the restriction and lifting maps
So defined, the conditional expectation $\varepsilon_{\mathbb{A}}$ is trace preserving, idempotent (so acts as a projector) and fulfills the bimodule property. The proofs of the latter are presented in Appendix A. The trace-preserving property of the conditional expectation $\varepsilon_{\mathbb{A}}$ can be easily verified as follows:
@@ Acknowledgements
Appendix A Idempotence and Bimodularity of $\varepsilon_{\mathbb{A}}$
In this appendix we prove two central properties of the conditional expectation $\varepsilon_{\mathbb{A}}$ defining the inclusion $\mathcal{T}\subset\mathcal{A}$: the idempotence, which amounts to say that $\varepsilon_{\mathbb{A}}$ is a projector from $\mathcal{A}$ to $\mathcal{T}$; and the bimodule property.
Idempotence.
We prove the idempotence of $\varepsilon_{\mathbb{A}}$ by computing $\varepsilon_{\mathbb{A}}\left(\tilde{\rho}\right)\in\mathcal{T}$ with $\tilde{\rho}=\alpha(\rho_{\mathbb{A}})$. Graphically this amounts to
$\rho$$\rho_{\mathbb{A}}$$\tilde{\rho}$${\varepsilon_{\mathbb{A}}\left(\tilde{\rho}\right)}$$\scriptstyle{\varepsilon_{\mathbb{A}}}$$\scriptstyle{\alpha}$$\scriptstyle{\varepsilon_{\mathbb{A}}}$
[…]
$\varepsilon_{\mathbb{A}}(\rho)$ $=\sum_{t\in\mathcal{T}}\,p_{t}\,\Pi_{t}\,,\quad p_{t}=\sum_{a\in\mathcal{A}}\,n^{t}_{a}\left(\frac{d_{t}}{d_{a}}\right)\,p_{a}\,$, (A.1)
and
$\tilde{\rho}$ $=\sum_{a\in\mathcal{A}}\,\widetilde{p}_{a}\,\Pi_{a}\,,\quad\widetilde{p}_{a}=\frac{1}{\lambda}\sum_{t\in\mathcal{T}}\,n^{t}_{a}\left(\frac{d_{a}}{d_{t}}\right)\,p_{t}\,.$ (A.2)
Therefore, we find that
[…]
The equality implying idempotence holds if $p_{t}=\tilde{p}_{t},\,\forall\,t\in\mathcal{T}$, which immediately follows from Eqs. (3.8a) and (3.8b).
Bimodule property.
We recall that this refers to
$\varepsilon_{\mathbb{A}}({\tilde{p}}\,{m}\,{\tilde{q}})={p}\,\varepsilon_{\mathbb{A}}({m})\,{q}\,\quad\forall\,{p},{q}\in\mathcal{T},{m}\in\mathcal{A},$
where the tildes refer to the lifted versions of operators in $\mathcal{T}$. To prove this property, we first focus on the first term where
${\tilde{p}}$ $=\sum_{b\in\mathcal{A}}\,{\tilde{p}}_{b}\,\Pi_{b}\,,\quad{\tilde{q}}=\sum_{c\in\mathcal{A}}\,{\tilde{q}}_{c}\,\Pi_{c}\,,\quad{m}=\sum_{d\in\mathcal{A}}\,{m}_{d}\,\Pi_{d}\,.$ (A.4)
Now, we note that for $a\in\mathcal{A}$,
$\mathbb{K}_{a}\,{\tilde{p}}$ $=\sum_{t\in\mathcal{T}}\,\sqrt{\lambda^{t}_{a}}\,\Lambda_{t,a}\left(\sum_{b\in\mathcal{A}}\,{\tilde{p}}_{b}\,\Pi_{b}\right)=\sum_{t\in\mathcal{T}}\,\sum_{b\in\mathcal{A}}\,\sqrt{\lambda^{t}_{a}}\,{\tilde{p}}_{b}\,\Lambda_{t,a}\,\Pi_{b}$ (A.5)
$=\sum_{t\in\mathcal{T}}\,\sqrt{\lambda^{t}_{a}}\,{\tilde{p}}_{a}\,\Lambda_{t,a}\,$,
where $\lambda^{t}_{a}=n^{t}_{a}\,\left(\frac{d_{t}}{d_{a}}\right)$. Similarly,
${\tilde{q}}\,\mathbb{K}^{\dagger}_{a}=\sum_{t\in\mathcal{T}}\,\sqrt{\lambda^{t}_{a}}\,{\tilde{q}}_{a}\,\Lambda^{\dagger}_{t,a}\,.$ (A.6)
With this,
[…]
$=\sum_{t\in\mathcal{T}}\,\left(\sum_{a\in\mathcal{A}}\,\lambda^{t}_{a}\,\left({\tilde{p}}_{a}{\tilde{q}}_{a}\,{m}_{a}\right)\right)\,\Pi_{t}=\sum_{t\in\mathcal{T}}\,c_{t}\,\Pi_{t}\,$,
with
$c_{t}=\sum_{a\in\mathcal{A}}\,\lambda^{t}_{a}\,\left({\tilde{p}}_{a}{\tilde{q}}_{a}\,{m}_{a}\right)\,.$ (A.8)
Now we focus on the second term where
${p}$ $=\sum_{r\in\mathcal{T}}\,{p}_{r}\,\Pi_{r}\,,\quad{q}=\sum_{s\in\mathcal{T}}\,{q}_{s}\,\Pi_{s}\,,\quad\varepsilon_{\mathbb{A}}({m})=\sum_{t\in\mathcal{T}}\,{m}_{t}\,\Pi_{t}\,$, (A.9)
in such a way that
${p}\,\varepsilon_{\mathbb{A}}({m})\,{q}$ $=\sum_{t,r,s\in\mathcal{T}}\,\left({p}_{r}\,{q}_{s}\,{m}_{t}\right)\,\delta_{rs}\delta_{st}\,\Pi_{t}$ (A.10)
$=\sum_{t\in\mathcal{T}}\,\left({p}_{t}\,{q}_{t}\,{m}_{t}\right)\,\Pi_{t}=\sum_{t\in\mathcal{T}}\,c_{t}\,\Pi_{t}\,$,
where the last equality uses Eqs. (3.18) and (A.8), thus proving the bimodule property of $\varepsilon_{\mathbb{A}}$.
Appendix B Dualities and equivalent condensations
@@
$\mathbb{K}^{\mathbb{A}}_{a}$ $=\sum_{t\in\mathcal{T}}\,\lambda_{ta}^{\mathbb{A}}\,\Lambda_{ta}\,,\quad\lambda_{ta}^{\mathbb{A}}=\sqrt{(n^{t}_{a})_{\mathbb{A}}\left(\frac{d_{t}}{d_{a}}\right)}\,$ (B.3)
@@
$\varepsilon_{\mathbb{A}}(\mathrm{Q}\,\rho\,\mathrm{Q}^{\dagger})\equiv\varepsilon_{\mathbb{B}}(\rho)$ (B.4)
@@
$\lambda_{ta}^{\mathbb{B}}=\sum_{b\in\mathcal{A}}\,q_{ab}\,\lambda_{tb}^{\mathbb{A}}$ (B.5)
@@
$\lambda_{\phi 1}^{\mathbb{A}}=\lambda_{\phi Y}^{\mathbb{A}}=1=\lambda_{\mathbb{X}X}^{\mathbb{A}}=\lambda_{\mathbb{X}Z}^{\mathbb{A}}=1\,$, (B.6)
@@
$\varepsilon_{\mathbb{A}}(\rho)$ $=p_{\phi}\,\Pi_{\phi}+p_{\mathbb{X}}\,\Pi_{\mathbb{X}}$ (B.7)
@@
$\lambda_{\phi 1}^{\mathbb{B}}=\lambda_{\phi Z}^{\mathbb{B}}=1=\lambda_{\mathbb{X}X}^{\mathbb{B}}=\lambda_{\mathbb{X}Y}^{\mathbb{B}}=1\,$, (B.8)
@@
$\varepsilon_{\mathbb{B}}(\rho)$ $=p_{\phi}\,\Pi_{\phi}+p_{\mathbb{X}}\,\Pi_{\mathbb{X}}$ (B.9)
@@
${\rm Q}=\left(\begin{array}[]{cc}\mathbb{I}_{2}&0\\ 0&X\end{array}\right)\,$, (B.10)
@@
$X=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\,$, (B.11)
@@
$\varepsilon_{\mathbb{A}}(\mathrm{Q}\,\rho\,\mathrm{Q}^{\dagger})\equiv\varepsilon_{\mathbb{B}}(\rho)\,.$ (B.12)